OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x,y). @ingenious help

OpenStudy (lalaly):

what are the equations

OpenStudy (anonymous):

x-2y=0 4x-3y=15

OpenStudy (lalaly):

solve for x in the first equation x=2y .... (1) now substitute (1) in the second equation 4(2y)-3y=15 8y-3y=15 5y=15 y=3 and we know that x=2y so x=2(3)= 6

OpenStudy (anonymous):

so itz (6,3)

OpenStudy (lalaly):

yup

OpenStudy (anonymous):

thx

OpenStudy (lalaly):

:)